OpenStudy (anonymous):

Use the properties of sigma notation and the summation formulas to evaluate the summation from i equals 1 to 10 of the quantity 2 times i squared plus 4 times i minus 5.

OpenStudy (anonymous):

https://gyazo.com/838e3b821bb0bb6c5de207edcd36eae4

OpenStudy (anonymous):

@jim_thompson5910 Would there be no shortcut? Other adding each one.

jimthompson5910 (jim_thompson5910):

you'll use these properties http://www.phengkimving.com/calc_of_one_real_var/09_the_intgrl/09_01_summ_notatn_and_formulas_files/image012.gif specifically property i and property ii

jimthompson5910 (jim_thompson5910):

and also \[\Large \sum_{i = 1}^{n}k = k*n\] where k is some constant

jimthompson5910 (jim_thompson5910):

\[\Large \sum_{i=1}^{10}\left(2i^2+4i-5\right) = \sum_{i=1}^{10}\left(2i^2\right)+\sum_{i=1}^{10}\left(4i\right)+\sum_{i=1}^{10}\left(-5\right)\] \[\Large \sum_{i=1}^{10}\left(2i^2+4i-5\right) = 2*\sum_{i=1}^{10}\left(i^2\right)+4*\sum_{i=1}^{10}\left(i\right)+\sum_{i=1}^{10}\left(-5\right)\] I'll let you finish

OpenStudy (anonymous):

I don't know what I did wrong I got 985

jimthompson5910 (jim_thompson5910):

Tell me what you got for \[\Large \sum_{i=1}^{10}\left(i^2\right)\] and \[\Large \sum_{i=1}^{10}\left(i\right)\]

OpenStudy (anonymous):

After multiplying by two I got 770 for the first one

OpenStudy (anonymous):

2nd one I got 220 after multiplying by 4

jimthompson5910 (jim_thompson5910):

\[\Large \sum_{i=1}^{10}\left(i^2\right)\] is not equal to 770

jimthompson5910 (jim_thompson5910):

oh you mean \[\Large 2*\sum_{i=1}^{10}\left(i^2\right)\] ???

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

\[\Large 2*\sum_{i=1}^{10}\left(i^2\right) = 770\] is correct

jimthompson5910 (jim_thompson5910):

\[\Large 4*\sum_{i=1}^{10}\left(i\right) = 4*55 = 220\] is also correct

jimthompson5910 (jim_thompson5910):

what is \[\Large \sum_{i=1}^{10}\left(-5\right)\] equal to?

OpenStudy (anonymous):

Won't it just be -5?

jimthompson5910 (jim_thompson5910):

nope

jimthompson5910 (jim_thompson5910):

think of it as you're summing 10 copies of (-5)

OpenStudy (anonymous):

-50

jimthompson5910 (jim_thompson5910):

smaller example \[\Large \sum_{i=1}^{4}\left(-5\right) = \left(-5\right)+\left(-5\right)+\left(-5\right)+\left(-5\right)\] notice how there are 4 copies of (-5) being added

jimthompson5910 (jim_thompson5910):

as a shortcut, you can say 10*(-5) = -50, so yes you have it

OpenStudy (anonymous):

Thank you good sir

OpenStudy (anonymous):

I have one final question I'd appreciate it if you could look over it.